In the coordinate plane connect points (1, 0) and (0, 2), (0, 2) and (2, 2), (2, 2) and (1, 4), (1, 4) and (5, 2), (5, 2) and (1, 0). What is the area of the polygon obtained?-3.5+ 2 15 -0.54 35 6n the coordinate plane connect points (1, 0) and (0, 2), (0, 2) and (2, 2), (2, 2) and (1, 4), (1, 4) and (5, 2), (5, 2) and (1, 0). What is the area of the polygon obtained?
Answer #1

10,10 -- 15,) (10) 2 3 Area of polynomial - Amo of A1+ A2 + 13. Ana of A = اردهال (K, J) - 11 x, y ) +4₂1 93-91) + x319-42) Afма Ръла 6 42 - - - ) + - 2-) - - 12(5) +р — 101 - з яе, ошь 43 = 12-(2-) + s(-5) + (-5). - ч+ (-10) — - 3 84 рр. — олса рвца
So here we divide the polygon in 3 triangles as shown in the figure and then we apply the triangle formula in coordinate geometry to find the area of 3 triangles and add them to find the area of the polynomial.